Session: Feb 2013 (session 1)
Wavelength: 1.3 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Roma at Wed Feb 27 17:06:27 2013) Success:
All scan recordeed. Sky clody. Not known problems.

Urumqi (Wenjun Yang at Thu Feb 28 03:17:11 2013) Success:

Torun (K. Borkowski at Thu Feb 28 06:42:21 2013) Success:
No known problems.

Hartebeesthoek (Jonathan Quick at Thu Feb 28 06:55:33 2013) Success:
No known problems, but pointing is probably poor

Shanghai (Bo Xia at Thu Feb 28 07:06:44 2013) Minor failures: RECO
We lost the first scan because of the recorder problem.

Yebes (Daniel Sastre at Thu Feb 28 17:08:48 2013) Success:
No known problems.

Onsala (Michael Lindqvist at Sat Mar 02 09:35:07 2013) Success:
Very foggy!

Metsahovi (E.Rastorgueva-Foi at Wed Mar 13 11:26:36 2013) Failure:
Both RCP and LCP (bbc's 01-04 and 05-08) were connected to the same IF of DBBC due to an unaccounted feature of Mh's DBBC construction.

Effelsberg (GEORGI at Fri Mar 15 08:08:54 2013) Success: